Dear colleagues,
Here is a summary of today's activites at CSBF.
- There was no power stoppage during the overnight muon run again. Larry told me that the heat blocking by the card board could reset the power supply. Also suggested to have the UPS connected to the GSE notebook computer so that the UPS power can be monitored. Or the Power GSE Labview program needs to have logging of the power data on. I will try to do both.
- Chris Field (CSBF) confirmed that CSBF power monitoring does not show any power off/on at that time. There are many CDM equipment with power on for 24 hours with power outlet next to what we are using and they did not have the problem.
- I confirmed with Dave that we could have the remote commanding and data flow this year since we still have the port waiver permission valid even though it has to be updated for another year asap.
- The CPT with CDM, which consists of 4 power modes, started from this morning.
- The Ether adapter cable, mating two 9 pin D-sub connectors with one 9 pin D-sub connector, was re-made with the mashed wire cables since the single wire cable is not qualified for flight and already failed in the acknowledgment connection.
- It took longer time than usual since not only we had SangEun as the main person but also it was the first CPT with the CDM for both Science and CDM.
- CDM side tests in muon run in flight mode took about 3 hours before the instrument test started.
- Common electronics tests from SFC1/2 shutdown-safe tests to the all detector data taking before each sub-detectors' LPT could be finished by 7:30 PM.
- It seems to me that this part of the test and together with the sub-detectors' LPTs (actually most of the instrument CPT) could be done only once during the CDM integration.
- The remaining LPT of sub-detectors for the first CPT will continue from tomorrow morning.
- For the next 3 CPTs with different power modes, will do mostly muon run in flight mode since only differences are in power mode not in any connections. Simulating the flight mode seems the best for the CDM test in the combined CPT.
- Manuals were updated based on today's experiences with changes in the cdaq software and parameters. Additional manual for the CDM integration CPT is being made.
- The repaired MXR box delivered and replaced with the PYL box which had a bad shaper channel. I thought it was a replacement of the PXR which is giving time out error. But the PXR box is a special box and Nick told me that it is not necessary to replace it with current error rate. The old PYL box will be shipped back to PSU for a repair. It will become a spare box.
- S3 threshold tuning continued. Nick clarified which values to be adjusted and definition of each channels as follows.
- SCI_TCD THRESHOLD, 8, 0x01, 500
- SCI_TCD THRESHOLD, 8, 0x05, 500
- (Change 500 to anything you wish > 100 and <=4095)
- The following should be helpful (but is true for S3 only):
- THR0 -> ZLo
- THR1 -> S3Lo
- THR2 -> Unused
- THR3 -> S3Hi
- Debugging of the reduced hsk data could not be tried today by Prashant due to the CPT.
- SangEun continue working on the SIP interface coding before and after the CPT work.
Here are things to do.
- Continuation of CPT with the CDM. Planning to have the UMD connection tomorrow with 1 hour commanding window in 2 hours and full data flow from 10 AM (EDT) to 5 PM (EDT). UMD SOC, please be prepared from 10 AM.
- S3 threshold value tuning.
- Reduced hsk debugging.
- Chercam HSK problem debugging.
- TCD 12 V HSK attenuation circuit installation.
- Update of manuals including CPT.
- Copying and backup of newly taken integration data into cdps3 at UMD and client PCs at CSBF.
- Adding another Ethernet card to cdps5 for a dual connection to local network and to outside.
Please let us know if you have any questions/comments/suggestions.
Cheers, MooHyun.
MooHyun, July 22, 2008



